OpenStudy (anonymous):

Solve for x and y. Show all work please. -3x+2y=2 7x-3y=7

OpenStudy (anonymous):

Does it want you to use elimination, substituion or does it not matter?

OpenStudy (anonymous):

Substitution Method

OpenStudy (anonymous):

If you're using substitution you have to get one on the equations to equal y so I'll choose the first one -3x+2y=2 Add 3x on both sides 2y=3x+2 Divide by 2 on both sides y=3/2x+1 Now take the equation above and substitute it into the second equation you were given So  Take out the y and put in (3/2x+1) and make sure it's in parentheses.

OpenStudy (anonymous):

So your equation will look like this after you substitute. 7x-3(3/2x+1)=7

OpenStudy (anonymous):

 That means multiply So 3*3/2x=? and 3*1=?